# Few Batches or Little Memory, But Not Both: Simultaneous Space and Adaptivity Constraints in Stochastic Bandits

### 14 Mar 2026 v1

@@ Abstract
We study stochastic multi-armed bandits under simultaneous constraints on space and adaptivity: the learner interacts with the environment in $B$ batches and has only $W$ bits of persistent memory. Prior work shows that each constraint alone is surprisingly mild: near-minimax regret $\widetilde{O}(\sqrt{KT})$ is achievable with $O(\log T)$ bits of memory under fully adaptive interaction, and with a $K$-independent $O(\log\log T)$-type number of batches when memory is unrestricted. We show that this picture breaks down in the simultaneously constrained regime. We prove that any algorithm with a $W$-bit memory constraint must use at least $\Omega(K/W)$ batches to achieve near-minimax regret $\widetilde{O}(\sqrt{KT})$ , even under adaptive grids. In particular, logarithmic memory rules out $K$-independent batch complexity.
Our proof is based on an information bottleneck. We show that near-minimax regret forces the learner to acquire $\Omega(K)$ bits of information about the hidden set of good arms under a suitable hard prior, whereas an algorithm with $B$ batches and $W$ bits of memory allows only $O(BW)$ bits of information. A key ingredient is a localized change-of-measure lemma that yields probability-level arm exploration guarantees, which is of independent interest. We also give an algorithm using $O(\log T)$ bits of memory and $\widetilde{O}(K)$ batches that achieves regret $\widetilde{O}(\sqrt{KT})$, which nearly matches our lower bound.
@@ 1 Introduction
Classical algorithms such as UCB (Auer et al., 2002) and Thompson sampling (Thompson, 1933; Agrawal and Goyal, 2012) achieve the minimax rate $\widetilde{O}(\sqrt{KT})$. These guarantees are typically established in an idealized model that implicitly grants two resources: enough memory to maintain rich per-arm statistics, and full adaptivity, meaning that the learner may update its policy after every observation.
@@
We model the space constraint by restricting the learner to $W$ bits of persistent memory, meaning that the state carried from one round to the next can take at most $2^{W}$ values. Computation within a round is unrestricted; only the state retained between rounds is limited. This is a natural model of space constrained computation. In bandit theory, it dates back to the finite-memory bandit formulations of Cover (1968); Cover and Hellman (2003), and is also consistent with the modern sublinear-space viewpoint of Liau et al. (2018). Under the standard word-RAM convention with word size $\Theta(\log T)$, $W=\Theta(\log T)$ corresponds to $O(1)$ machine words.
@@
We consider the batched bandit model (Perchet et al., 2016; Gao et al., 2019), in which the horizon is partitioned into $B$ batches and outcomes are only incorporated between batches. At the beginning of each batch, the learner must commit to the action-generation rule for that batch using only information available at the previous batch boundary. Equivalently, feedback observed within a batch cannot affect later actions in the same batch. Following Gao et al. (2019), we consider both the static-grid and the adaptive-grid variants. The static grid assumes that the batch boundaries are predetermined and fixed in advance, while the adaptive grid allows the learner to choose the next batch boundary using information available at the previous boundary.
Viewed separately, each constraint appears surprisingly mild. Stochastic bandits admit near-optimal algorithms using only $O(\log T)$ bits of memory under fully adaptive interaction (Cover and Hellman, 2003; Liau et al., 2018), while batched bandits attain near-minimax regret with only $O(\log\log T)$ batches when memory is unrestricted (Perchet et al., 2016; Gao et al., 2019; Jin et al., 2021b; Esfandiari et al., 2021). These positive results rely on the other resource being abundant: space-bounded algorithms exploit frequent updates, whereas few-batch algorithms retain rich statistics across batch boundaries.
We consider the joint regime where the learner has a $W$-bit memory budget and interacts with the environment in $B$ batches. Our results show that this regime is qualitatively different from either constraint in isolation: once the space budget is logarithmic, near-minimax regret is no longer compatible with a $K$-independent number of batches.
@@
We also show that this lower bound is essentially tight in the natural logarithmic-memory regime.
@@
In the regime $W=\Theta(\log T)$, there exists a stochastic static-grid bandit algorithm using $O(\log T)$ bits of memory and $\widetilde{O}(K)$ batches that achieves near-minimax regret $\widetilde{O}(\sqrt{KT}).$
A formal version of Theorem 2 is given in Section˜5.
@@
  • Two technical ingredients for the lower bound. We establish (i) a probability-level per-good-arm exploration lemma, and (ii) a localized change-of-measure lemma for under-sampling events, which lets the proof pay only for the first $n$ pulls of the perturbed arm.
  • A near-matching upper bound in the logarithmic-memory regime. In the regime $W=\Theta(\log T)$, we provide a static-grid algorithm using $O(\log T)$ bits of memory and $\widetilde{O}(K)$ batches that achieves regret $\widetilde{O}(\sqrt{KT})$.
@@
It is obvious that the adaptive-grid model has stronger algorithmic power than the static-grid model, because the learner can choose the batch boundaries adaptively. Our lower bound is proved for the algorithmically stronger adaptive-grid model, while our upper bound is achieved in the algorithmically weaker static-grid model. Therefore, in each setting, our result is established in the strongest possible sense.
@@ 1.1 Technical Overview
$Y_{i}:=\mathbb{I}\{N_{i}(T)\geq n\},\qquad Y\in\{0,1\}^{K}.$
Thus $Y_{i}$ records whether arm $i$ received a nontrivial share of the exploration budget.
@@
What is missing is the complementary statement that genuinely good arms must also cross the threshold with constant probability. This is the delicate part of the proof. Let $\nu_{S^{\star}}$ be the realized instance under this prior, and on that instance alone one might imagine that it suffices to find some good arm and then exploit it. However, we prove that we need to find every good arm with constant probability. The key point is that the regret guarantee is uniform over all instances, not tailored to a single $\nu_{S}$. For every $j\in S^{\star}$, we consider a neighboring instance $\nu_{S^{\star}}^{(j)}$ in which arm $j$ has mean $1/2+\Delta_{0}$, for some perturbation size $\Delta_{0}>0$ to be specified later, while all other arm means remain the same as in $\nu_{S^{\star}}$. Thus arm $j$ becomes uniquely optimal under $\nu_{S^{\star}}^{(j)}$. If the learner largely ignores arm $j$ under the base instance, then it is poorly positioned for $\nu_{S^{\star}}^{(j)}$ and would incur large regret there. Hence no genuinely good arm can be safely ignored.
@@
To make the previous intuition formal, we need a probability-level statement: for every good arm $j$,
$\mathbb{P}_{\nu_{S}}(N_{j}(T)\geq n)\geq\beta$
for some constant $\beta>0$. An expectation lower bound on $N_{j}(T)$ is too weak for this purpose. Indeed, a learner could ignore arm $j$ on most trajectories and explore it extremely heavily only on a tiny exceptional set of runs. Such bursty exploration can keep $\mathbb{E}[N_{j}(T)]$ at scale $n$ while making $\mathbb{P}(N_{j}(T)\geq n)$ arbitrarily small. On the base instance $\nu_{S}$, this pathology is not ruled out by regret, because replacing pulls of one good arm by pulls of other good arms incurs no additional regret.
A natural route is therefore to argue that, on the perturbed instance $\nu_{S}^{(j)}$, near-minimax regret forces arm $j$ to be sampled at least $n$ times with sufficiently large probability, and then transfer this statement back to the base instance $\nu_{S}$. The obstacle is that standard transcript-level change-of-measure bounds are poorly aligned with the under-sampling event
$E_{j}:=\{N_{j}(T)<n\}.$
Standard change-of-measure bounds compare the full laws of the two experiments and therefore pay for all pulls of arm $j$, whereas the event $E_{j}$ depends only on whether the first $n$ pulls of arm $j$ ever occur. A useful toy picture is the following: under the perturbed instance, the learner may use the first $n$ pulls of arm $j$ to discover that $j$ is special and then exploit $j$ for the remaining TnT-n rounds. The event $E_{j}$ is already decided after those first $n$ pulls, so all later exploitation pulls are irrelevant to $E_{j}$. A standard full-transcript comparison nevertheless pays for all pulls of arm $j$, including those post-identification exploitation pulls, and can therefore incur a cost as large as order $T\Delta_{0}^{2}$, which is far too large for our parameter regime.
Our localized change-of-measure lemma truncates this comparison to the first $n$ pulls relevant to $E_{j}$, so the transfer cost is only of order $n\Delta_{0}^{2}$. With the perturbation gap chosen at the nearly minimax-detectable scale
@@
This is exactly what allows the desired probability-level exploration guarantee to be transferred back from $\nu_{S}^{(j)}$ to $\nu_{S}$.
@@
The upper bound in the logarithmic-memory regime matches the same bottleneck from the algorithmic side. Instead of storing a large active set of candidate arms, the algorithm keeps only an incumbent arm and a small summary of its performance, and then scans the remaining arms one by one as challengers. Each challenger is compared with the incumbent through a multistage test. This sequentialization deliberately trades memory for adaptivity: the memory remains logarithmic, while the number of batches becomes nearly linear in $K$. The resulting algorithm uses $O(\log T)$ bits of memory and $\widetilde{O}(K)$ batches, nearly matching the lower bound up to polylogarithmic factors.
@@ 2 Related Work
A large literature studies bandits under restrictions on how often the learner may update its policy. For stochastic batched bandits, early results include Perchet et al. (2016) for the two-armed case and Gao et al. (2019), which developed the BaSE policy and established near-optimal guarantees and lower bounds for general $K$-armed problems. Subsequent work extended these ideas to broader settings, including stochastic, linear, and adversarial bandits (Esfandiari et al., 2021), Thompson-sampling-based batched methods (Kalkanli and Ozgur, 2021), and anytime batched algorithms with near-optimal $O(\log\log T)$-type batch complexity (Jin et al., 2021b). Related guarantees are also known in structured settings such as finite-action linear contextual bandits (Han et al., 2020; Ruan et al., 2021), high-dimensional sparse linear contextual bandits (Ren and Zhou, 2024), generalized linear contextual bandits (Ren et al., 2020; Sawarni et al., 2024), and large-action-space contextual linear bandits (Hanna et al., 2023; Ren et al., 2024). Related formulations restrict adaptivity through switching costs, delayed feedback, or more general round-complexity constraints (Arora et al., 2012; Cesa-Bianchi et al., 2013; Joulani et al., 2013; Vernade et al., 2020; Agarwal et al., 2017; Jun et al., 2016). However, these works typically bound how often the policy may change while allowing the learner to retain arbitrarily rich information between updates.
@@
Learning with limited memory has been studied broadly in learning theory (Raz, 2016; Garg et al., 2017; Moshkovitz and Moshkovitz, 2017; Moshkovitz and Tishby, 2017). Within bandits, classical antecedents on finite-memory models go back to Cover (1968); Cover and Hellman (2003), and modern stochastic bandit results show that one can achieve strong regret guarantees with constant or logarithmic space under fully sequential interaction (Liau et al., 2018). Related memory-limited variants include memory-constrained adversarial bandits (Xu and Zhao, 2021), limited-memory approaches based on subsampling or forgetting (Baudry et al., 2021), and bounded arm-memory models in which the learner may keep statistics for only a limited number of arms at a time (Chaudhuri and Kalyanakrishnan, 2020; Maiti et al., 2021).
A different notion appears in streaming bandits, where arms arrive in a stream and the learner is constrained by pass and storage limitations (Assadi and Wang, 2020; Jin et al., 2021a; Agarwal et al., 2022; Assadi and Wang, 2022; Wang, 2023; Li et al., 2023; Assadi and Wang, 2024). These models are not directly comparable to ours. In the streaming formulations above, the available arm set changes over time, whereas in our setting the arm set is fixed and always accessible. Moreover, streaming algorithms are typically fully adaptive within a pass, while in our batched model the learner cannot use feedback observed within a batch to change later actions in that batch.
@@ 4 Lower Bounds on Batch Complexity under Space Constraints
If $W=O(\log T)$, then
$B=\Omega(K/\log T).$
The regret requirement in Theorem 1 is chosen to exactly match the known batched-bandit upper bound (Gao et al., 2019)
$O(\sqrt{KT}\log T\log K).$
The theorem shows that even attaining this rate under a $W$-bit persistent-memory budget requires batch complexity at least on the order of $K/W$.
@@ 4.1 An Event-Restricted Change of Measure and Per-Good-Arm Exploration
At the same time, the underlying change-of-measure argument is not specific to our Bernoulli hard family. For independent interest, we therefore state it in a general adaptive-sampling framework in which two environments differ only through a single data stream. This requires us to make the filtration and measurability structure explicit. In the present paper, however, no additional measurability difficulty arises: in our Bernoulli bandit application, all such properties are automatic under the standard construction of the interaction process. The extra generality is included only to isolate a reusable tool.

### 30 Mar 2026 v2

@@ Abstract
We study stochastic multi-armed bandits under simultaneous constraints on space and adaptivity: the learner interacts with the environment in $B$ batches and has only $W$ bits of persistent memory. Prior work shows that each constraint alone is surprisingly mild: near-minimax regret $\widetilde{O}(\sqrt{KT})$ is achievable with $O(\log T)$ bits of memory under fully adaptive interaction, and with a $K$-independent $O(\log\log T)$-type number of batches when memory is unrestricted. We show that this picture breaks down in the simultaneously constrained regime. We prove that any algorithm with a $W$-bit memory constraint must use at least $\Omega(K/W)$ batches to achieve near-minimax regret $\widetilde{O}(\sqrt{KT})$, even under adaptive grids. In particular, logarithmic memory rules out $O(K^{1-\varepsilon})$ batch complexity.
Our proof is based on an information bottleneck. We show that near-minimax regret forces the learner to acquire $\Omega(K)$ bits of information about the hidden set of good arms under a suitable hard prior, whereas an algorithm with $B$ batches and $W$ bits of memory allows only $O(BW)$ bits of information. A key ingredient is a localized change-of-measure lemma that yields probability-level arm exploration guarantees, which is of independent interest. We also give an algorithm that, for any bit budget $W$ with $\Omega(\log T)\leq W\leq O(K\log T)$, uses at most $W$ bits of memory and $\widetilde{O}(K/W)$ batches while achieving regret $\widetilde{O}(\sqrt{KT})$, nearly matching our lower bound up to polylogarithmic factors.
footnotetext: Equal contribution. Corresponding author.
@@ 1 Introduction
Classical algorithms such as UCB [ACF02] and Thompson sampling [THO33, AG12] achieve the minimax rate $\widetilde{O}(\sqrt{KT})$. These guarantees are typically established in an idealized model that implicitly grants two resources: enough memory to maintain rich per-arm statistics, and full adaptivity, meaning that the learner may update its policy after every observation.
@@
We model the space constraint by restricting the learner to $W$ bits of persistent memory, meaning that the state carried from one round to the next can take at most $2^{W}$ values. Computation within a round is unrestricted; only the state retained between rounds is limited. This is a natural model of space-constrained computation. In bandit theory, it dates back to the finite-memory bandit formulations of [COV68, CH03], and is also consistent with the modern sublinear-space viewpoint of [LSP+18]. Under the standard word-RAM convention with word size $\Theta(\log T)$, $W=\Theta(\log T)$ corresponds to $O(1)$ machine words.
@@
We consider the batched bandit model [PRC+16, GHR+19], in which the horizon is partitioned into $B$ batches and outcomes are only incorporated between batches. At the beginning of each batch, the learner must commit to the action-generation rule for that batch using only information available at the previous batch boundary. Equivalently, feedback observed within a batch cannot affect later actions in the same batch. Following [GHR+19], we consider both the static-grid and the adaptive-grid variants. The static grid assumes that the batch boundaries are predetermined and fixed in advance, while the adaptive grid allows the learner to choose the next batch boundary using information available at the previous boundary.
Viewed separately, each constraint appears surprisingly mild. Stochastic bandits admit near-optimal algorithms using only $O(\log T)$ bits of memory under fully adaptive interaction [CH03, LSP+18], while batched bandits attain near-minimax regret with only $O(\log\log T)$ batches when memory is unrestricted [PRC+16, GHR+19, JTX+21b, EKM+21]. These positive results rely on the other resource being abundant: space-bounded algorithms exploit frequent updates, whereas few-batch algorithms retain rich statistics across batch boundaries.
We consider the joint regime where the learner has a $W$-bit memory budget and interacts with the environment in $B$ batches. Our results show that this regime is qualitatively different from either constraint in isolation: once the space budget is logarithmic, near-minimax regret is no longer compatible with an $O(K^{1-\varepsilon})$ number of batches.
@@
We also show that this lower bound is essentially tight.
@@
For every bit budget $\Omega(\log T)\leq W\leq O(K\log T)$ , there exists a stochastic static-grid bandit algorithm using at most $W$ bits of persistent memory and $\widetilde{O}(K/W)$ batches that achieves near-minimax regret $\widetilde{O}(\sqrt{KT})$.
A formal version of Theorem 2 is given in Section˜5. Here the $\Omega(\log T)\leq W$ lower bound is needed to represent the relevant counters and empirical means to inverse-polynomial precision in $T$, while the $W\leq O(K\log T)$ upper bound is without loss of generality since standard finite-armed bandit algorithms only need to store per-arm summaries for the $K$ arms.
@@
  • Two technical ingredients of independent interest. In proving our lower bound, we establish two technical lemmas of independent interest: (i) a probability-level per-good-arm exploration lemma, and (ii) a localized change-of-measure lemma for under-sampling events, which lets the proof pay only for the first $n$ pulls of the perturbed arm.
  • A near-matching upper bound across memory budgets. We propose a static-grid algorithm such that, for any bit budget $\Omega(\log T)\leq W\leq O(K\log T)$, it uses at most $W$ bits of persistent memory, at most $\widetilde{O}(K/W)$ batches, and achieves near-minimax regret $\widetilde{O}(\sqrt{KT})$.
@@
The adaptive-grid model has stronger algorithmic power than the static-grid model, because the learner can choose the batch boundaries adaptively. Our lower bound is proved for the algorithmically stronger adaptive-grid model, while our upper bound is achieved in the algorithmically weaker static-grid model. Therefore, in each setting, our result is established in the strongest possible sense.
@@ 1.1 Technical Overview
$Y_{i}:=\mathbb{I}\{N_{i}(T)\geq n\},\qquad Y\in\{0,1\}^{K}$
where $N_{i}(T)$ means the number of pulls of arm $i$ up to the horizon $T$. Thus $Y_{i}$ records whether arm $i$ received a nontrivial share of the exploration budget.
@@
What is missing is the complementary statement that genuinely good arms must also cross the threshold with constant probability. This is the delicate part of the proof. Let $\nu_{S^{\star}}$ be the realized instance under this prior, and on that instance alone one might imagine that it suffices to find some good arm and then exploit it. However, we prove that every good arm must cross the threshold with constant marginal probability. The key point is that the regret guarantee is uniform over all instances, not tailored to a single $\nu_{S}$. For every $j\in S^{\star}$, we consider a neighboring instance $\nu_{S^{\star}}^{(j)}$ in which arm $j$ has mean $1/2+\Delta_{0}$, for some perturbation size $\Delta_{0}>0$ to be specified later, while all other arm means remain the same as in $\nu_{S^{\star}}$. Thus arm $j$ becomes uniquely optimal under $\nu_{S^{\star}}^{(j)}$. If the learner largely ignores arm $j$ under the base instance, then it is poorly positioned for $\nu_{S^{\star}}^{(j)}$ and would incur large regret there. Hence no genuinely good arm can be safely ignored.
@@
To make the previous intuition formal, we need a probability-level statement: for every good arm $i$,
$\mathbb{P}_{\nu_{S}}(N_{i}(T)\geq n)\geq\beta$
for some constant $\beta>0$. An expectation lower bound on $N_{i}(T)$ is too weak for this purpose. Indeed, a learner could ignore arm $i$ on most trajectories and explore it extremely heavily only on a tiny exceptional set of runs. Such bursty exploration can keep $\mathbb{E}[N_{i}(T)]$ at scale $n$ while making $\mathbb{P}(N_{i}(T)\geq n)$ arbitrarily small. On the base instance $\nu_{S}$, this pathology is not ruled out by regret, because replacing pulls of one good arm by pulls of other good arms incurs no additional regret.
A natural route is therefore to argue that, on the perturbed instance $\nu_{S}^{(i)}$, near-minimax regret forces arm $i$ to be sampled at least $n$ times with sufficiently large probability, and then transfer this statement back to the base instance $\nu_{S}$. The obstacle is that standard transcript-level change-of-measure bounds are poorly aligned with the under-sampling event
$E_{i}:=\{N_{i}(T)<n\}.$
Standard change-of-measure bounds compare the full laws of the two experiments and therefore pay for all pulls of arm $i$, whereas the event $E_{i}$ depends only on whether the first $n$ pulls of arm $i$ ever occur. A useful toy picture is the following: under the perturbed instance, the learner may use the first $n$ pulls of arm $i$ to discover that $i$ is special and then exploit $i$ for the remaining TnT-n rounds. The event $E_{i}$ is already decided after those first $n$ pulls, so all later exploitation pulls are irrelevant to $E_{i}$. A standard full-transcript comparison nevertheless pays for all pulls of arm $i$, including those post-identification exploitation pulls, and can therefore incur a cost as large as order $T\Delta_{0}^{2}$, which is far too large for our parameter regime.
Our localized change-of-measure lemma truncates this comparison to the first $n$ pulls relevant to $E_{i}$, so the transfer cost is only of order $n\Delta_{0}^{2}$. With the perturbation gap chosen at the nearly minimax-detectable scale
@@
This is exactly what allows the desired probability-level exploration guarantee to be transferred back from $\nu_{S}^{(i)}$ to $\nu_{S}$.
@@
The upper bound matches the same bottleneck from the algorithmic side. Instead of storing a large active set of candidate arms, the algorithm keeps only an incumbent arm, a small summary of its performance, and a block of at most $S$ challenger arms at a time. These challengers are compared with the incumbent through a multistage test, after which the algorithm moves on to the next block. This blockwise scan explicitly trades memory for adaptivity: with block size $S$, the implementation uses $O(S\log T)$ bits of memory and the number of batches drops to $\widetilde{O}(K/S)$. Equivalently, for any bit budget $W$ with $\Omega(\log T)\leq W\leq O(K\log T)$, choosing $S\asymp W/\log T$ gives an implementation using at most $W$ bits and $\widetilde{O}(K/W)$ batches. The resulting algorithm nearly matches the lower bound up to polylogarithmic factors.
Note that the $W\leq O(K\log T)$ upper bound is without loss of generality since standard finite-armed bandit algorithms only need to store per-arm summaries for the $K$ arms.
@@ 2 Related Work
A large literature studies bandits under restrictions on how often the learner may update its policy. For stochastic batched bandits, early results include [PRC+16] for the two-armed case and [GHR+19], which developed the BaSE policy and established near-optimal guarantees and lower bounds for general $K$-armed problems. Subsequent work extended these ideas to broader settings, including stochastic, linear, and adversarial bandits [EKM+21], Thompson-sampling-based batched methods [KO21], and anytime batched algorithms with near-optimal $O(\log\log T)$-type batch complexity [JTX+21b]. Related guarantees are also known in structured settings such as finite-action linear contextual bandits [HZZ+20, RYZ21], high-dimensional sparse linear contextual bandits [RZ24], generalized linear contextual bandits [RZK20, SDB+24], and large-action-space contextual linear bandits [HYF23, RJX24]. Related formulations restrict adaptivity through switching costs, delayed feedback, or more general round-complexity constraints [ADT12, CDS13, JGS13, VCL+20, AAA+17, JJN+16]. However, these works typically bound how often the policy may change while allowing the learner to retain arbitrarily rich information between updates.
@@
Learning with limited memory has been studied broadly in learning theory [RAZ16, GRT17, MM17, MT17]. Within bandits, classical antecedents on finite-memory models go back to [COV68, CH03], and modern stochastic bandit results show that one can achieve strong regret guarantees with constant or logarithmic space under fully sequential interaction [LSP+18]. Related memory-limited variants include memory-constrained adversarial bandits [XZ21], limited-memory approaches based on subsampling or forgetting [BRC21], and bounded arm-memory models in which the learner may keep statistics for only a limited number of arms at a time [CK20, MPK21].
A different notion appears in streaming bandits, where arms arrive in a stream and the learner is constrained by pass and storage limitations [AW20, JHT+21a, AKP22, AW22, WAN23, LZW+23, AW24]. These models are quite different from ours. In the streaming formulations above, the available arm set changes over time, whereas in our setting the arm set is fixed and always accessible. Moreover, streaming algorithms are typically fully adaptive within a pass, while in our batched model the learner cannot use feedback observed within a batch to change later actions in that batch. This distinction is also reflected in the complexity trade-off: in streaming bandits, one can recover the optimal regret rate up to logarithmic factors with constant space and only $O(\log\log T)$ passes, whereas in our batched model, recovering the same rate with $O(\log T)$ bits of memory requires $\widetilde{\Omega}(K)$ batches. Furthermore, techniques used in proving lower bounds are fundamentally different between streaming bandits and our setting.
@@ 4 Lower Bounds on Batch Complexity under Space Constraints
Theorem˜1 applies to any $\tilde{O}(\sqrt{KT})$ regret bound with minor modifications; we adopt the concrete rate $O(\sqrt{KT}\log T\log K)$ here to exactly match the known batched-bandit upper bound [GHR+19]. Thus, the theorem shows that even attaining this rate under a $W$-bit persistent-memory budget requires batch complexity at least on the order of $K/W$.
@@ 4.1 An Event-Restricted Change of Measure and Per-Good-Arm Exploration
At the same time, the underlying change-of-measure argument is not specific to our Bernoulli hard family. For independent interest, we state it in a reusable adaptive-sampling form. To keep the main proof focused on the lower-bound argument, the full measurable-space formulation and the associated truncation-based proofs are deferred to Appendix A.1.